OpenStudy (cheyfae7):

which is equal to 8^-10/8^5 ? a. 8^50 b. 8^-15 c. 8^-5 d. 8^-2

OpenStudy (alexandervonhumboldt2):

\[\frac{ a^b }{ a^c }=a^{b-c}\]

OpenStudy (alexandervonhumboldt2):

thus \[\frac{ 8^{-10} }{ 8^5 }=8^{-10-5}=?\]

OpenStudy (cheyfae7):

thats what i thought so the answer would be c . 8^-5

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (cheyfae7):

aahhh i see now 8^-15

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (cheyfae7):

thanks again !

OpenStudy (alexandervonhumboldt2):

:)